OpenStudy (anonymous):

"The start up cost of a company is 150,000$. It costs $.17 to manufacture each headset/ graph the function that represents the average cost of a headset. how many must be manufactured to result in a cost less then 5$ per headset." If someone could please help, that'd be great. I don't need helping graphing, but if you could show me how to set up the equation, I'd love it! Thanks. :)

OpenStudy (anonymous):

@phi, can you please help?

OpenStudy (phi):

one way to think about it: i f you made one headset, it would cost 150,000 (the start-up cost) and then another 0.17 to make the headset. if you made 2 headsets, it still costs 150,000 for start-up, and 2*0.17 to make the two. the total cost to make the 2 headsets is T= 150,000 + 2*0.71 the average cost (or cost per headset) is T/2 can you use these ideas to figure out the average cost for N headsets ?

OpenStudy (phi):

you want to find an equation to answer how many must be manufactured to result in a cost less then 5$ per headset." it would look something like T/N < 5 where T is the total cost to make N headsets, and N is the number of headsets.

OpenStudy (anonymous):

The cost is to start with (as you have spent 150,000 that will give you a minus sign) \[m=-150,000\] Production cost p: \[p=-0.17x\] And the cost to consumer is: \[h=5x\] where x is number of units. The linear function for this companys economy would then be something like... y= PROFIT-COST Did it help? :)

OpenStudy (anonymous):

@phi, where did you get 2 in the average cost? = T/2??

OpenStudy (phi):

that was an example for just 2 headsets you have to extend the idea to N head sets

OpenStudy (anonymous):

Also, @kalph , you said y=p-c but which cost specifically is being subtracted? You mentioned 3...

OpenStudy (anonymous):

Okay, so then my equation should look like: 5=150,000 +1.7x and we solve for x. Does that look right @Phi?

OpenStudy (phi):

no. what are the chances 5 = 150,000 + more stuff ? The idea is write down how much it costs to make N headsets. The way to figure that out is write down how much it costs to make 1 headset, or 2 or 3... and llook for the pattern. that is the first step.

OpenStudy (anonymous):

okay, so T/n<5 t/1<5, n =1; t=5 t/2<5, n=2; t=10 t/3 <5, n=3; t=15 So we can conclude that it varies directly, right? and that's a step in the right direction?

OpenStudy (phi):

the cost is the startup cost plus the manufacturing cost. the cost to build 1 headset is a lot: 150,000.17

OpenStudy (anonymous):

So we know the equation is y=kx or y/x=k y=t=150,000 (total cost available?) x=n.17=Cost per set being manufactured. This fraction when simplified should equal 5, right? 150,000/n.17 =5

OpenStudy (anonymous):

why is the cost to build a headset 150,000.17?? That exceeds the budget- or am I misunderstanding something?

OpenStudy (phi):

"The start up cost of a company is 150,000$. It costs $.17 to manufacture each headset/ graph the function that represents the average cost of a headset. You include the cost to start-up.

OpenStudy (anonymous):

okay. Thanks.

OpenStudy (anonymous):

Nothing i'm setting up is working out. Can you please give me the equation and what each variable is so I can solve? I'd really appreciate it. :)

OpenStudy (anonymous):

@phi?? Hello??

OpenStudy (phi):

what is the cost if you made *only* 3 headsets, including start up costs?

OpenStudy (anonymous):

150,000.51

OpenStudy (phi):

now if it were N headsets, what would you write down ?

OpenStudy (anonymous):

150,000 + n(.17) right? @phi But what does it equal?? Like you said, it can't equal 5.

OpenStudy (phi):

that is the cost to make n headsets how many must be manufactured to result in a cost less then 5$ per headset. ? what is the total cost of n headsets if each cost $5 ? answer: 5n you want the total cost to be less than 5n

OpenStudy (phi):

150,000 + n(.17) < 5n that is now quite the function they are asking for. graph the function that represents the average cost of a headset if the total cost is 150,000 + n(.17) what is the average cost? the average is the total divided by n so they want you to graph y= (150,000 + x(.17) )/x= 150,000/x + 0.17 (I changed n to x, so you can plot it on an x-y graph)

OpenStudy (phi):

It should be obvious that it will take a lot of headsets to get the average cost down to $5.00

OpenStudy (phi):

if they cost nothing to make (meaning ignore the 0.17) then we could divide $5 into 150,000 to get 30,000 headsets It will take a bit more than that because they do cost a little bit to make each one. but you can solve 150,000 + n(.17) < 5n add -0.17n to both sides 150,000 < 5n - 0.17n 150,000 < (5-0.17) n (factor out the n 150,000 < 4.83 n (simplify 5-0.17) divide both sides by 4.83 150,000/4.83 < n 31055.9006 < n that means n> 31056

OpenStudy (anonymous):

While you were typing that out, I was solving and got the same thing. I know this was hard cause I wasn't exactly getting it, but I do now!! Thank You SO much!! :D

OpenStudy (anonymous):

Can you just read through a word problem for me and tell me if it's a trick question?? I already have the answer, but I'd love if you'd just check my work. :) @phi

OpenStudy (phi):

did you post it?